### v1

# Coexistence and Extinction in Flow-Kick Systems: An invasion growth rate approach

Abstract. Natural populations experience a complex interplay of continuous and discrete processes: continuous growth and interactions are punctuated by discrete reproduction events, dispersal, and external disturbances. These dynamics can be modeled by impulsive or flow-kick systems, where continuous flows alternate with instantaneous discrete changes. To study species persistence in these systems, an invasion growth rate theory is developed for flow-kick models with state-dependent timing of kicks and auxiliary variables that can represent stage structure, trait evolution, and environmental forcing. The invasion growth rates correspond to Lyapunov exponents that characterize the average per-capita growth of species when rare. Two theorems are proven that use invasion growth rates to characterize permanence, a form of robust coexistence where populations remain bounded away from extinction. The first theorem uses Morse decompositions of the extinction set and requires that there exists a species with a positive invasion growth rate for every invariant measure supported on a component of the Morse decomposition. The second theorem uses invasion growth rates to define invasion graphs whose vertices correspond to communities and directed edges to potential invasions. Provided the invasion graph is acyclic, permanence and extinction are fully characterized by the signs of the invasion growth rates. Invasion growth rates are also used to identify the existence of extinction-bound trajectories and attractors that lie on the extinction set. To demonstrate the framework’s utility, these results are applied to three ecological systems: (i) a microbial serial transfer model where state-dependent timing enables coexistence through a storage effect, (ii) a spatially structured consumer-resource model showing intermediate reproductive delays can maximize persistence, and (iii) an empirically parameterized Lotka-Volterra model demonstrating how disturbance can lead to extinction by disrupting facilitation. Mathematical challenges, particularly for systems with cyclic invasion graphs, and promising biological applications are discussed. These results reveal how the interplay between continuous and discrete dynamics creates ecological outcomes not found in purely continuous or discrete systems, providing a foundation for predicting population persistence and species coexistence in natural communities subject to gradual and sudden changes.Keywords: impulsive differential equations, flow-kick systems, uniform persistence, permanence, Lyapunov exponents, growth rates when rare, coexistence, extinction.
1. Introduction
Many biological systems naturally combine continuous and discrete processes. For example, animal populations may grow and interact continuously throughout a season but reproduce only at specific times [Gyllenberg et al., 1997, Pachepsky et al., 2008, Geng and Lutscher, 2021, Geng et al., 2021], diseases may spread continuously while vaccination programs occur in discrete pulses [Agur et al., 1993, Meng and Chen, 2008, d’Onofrio, 2002, Bai, 2015], lakes may experience continuous nutrient cycling interrupted by sudden influxes from storm events [Meyer et al., 2018], and microbial communities in serial-transfer experiments alternate between continuous growth and dilution events [Yi and Dean, 2013, Wolfe and Dutton, 2015, Venkataram et al., 2016, Yurtsev et al., 2016, Good et al., 2017, Letten and Ludington, 2023]. Traditional purely continuous or discrete models fail to capture these hybrid dynamics, potentially missing important emergent properties and complex behaviors that arise from the interaction between continuous and discrete processes.
To address this reality, impulsive differential equations (also known as flow-kick models) combine continuous-time and discrete-time processes [Lakshmikantham et al., 1989, Meyer et al., 2018]. In these models, continuous processes are modeled by the flow of a system of ordinary differential equations (ODEs). Instantaneous changes in the system (the “kicks”) occur at specific times and are modeled by an update map. The complete model alternates between the flow and the kicks: the system flows according to its ODEs for a set period of time $\tau$, experiences instantaneous kicks, then flows again, establishing a recurring pattern of disturbance and recovery. The time $\tau$ between kicks can be fixed or state-dependent, depending on the biological context. In models of seasonal reproduction, for example, kicks occur at fixed annual intervals. In contrast, the timing of kicks may depend on system states, as seen in disease models where vaccination pulses are triggered by disease prevalence [Agur et al., 1993, d’Onofrio, 2002, Meng and Chen, 2008, Bai, 2015], or in serial transfer experiments where dilution events are determined by the optical density of the medium [Yi and Dean, 2013]. The kicks can correspond to changes in population state (e.g. a dilution event in a serial transfer experiment or a reproductive pulse) or changes in flow (e.g. seasonal changes in demographic rates).
Flow-kick systems present unique mathematical challenges for examining questions of coexistence and extinction. A fruitful approach in theoretical population biology to study these questions uses invasion growth rates—the average per-capita growth rate of a species or genotype when rare [MacArthur and Levins, 1967, Chesson, 1982, 1994, Ellner et al., 2016, Barabás et al., 2018, Ellner et al., 2019, Spaak and Schreiber, 2023, Clark et al., 2024]. Mathematically, these invasion growth rates correspond to Lyapunov exponents (a measure of how perturbations grow or shrink over time) and can be used to characterize a strong form of coexistence called permanence [Schreiber, 2000, Schreiber et al., 2011, Hofbauer and Schreiber, 2022]. Permanence occurs when a system maintains a global attractor bounded away from extinction [Schuster et al., 1979, Sigmund and Schuster, 1984, Butler et al., 1986], ensuring that coexistence remains robust to both small, repeated perturbations and rare, large disturbances [Jansen and Sigmund, 1998, Schreiber, 2006]. Invasion growth rates also provide an approach to identify the existence of extinction-bound trajectories – trajectories where all species are initially present but where a subset of species ultimately tends toward extinction. Although the invasion growth rate approaches to permanence and extinction are well developed for purely continuous or discrete-time models [Rand et al., 1994, Ashwin et al., 1994, Garay and Hofbauer, 2003, Roth et al., 2017, Patel and Schreiber, 2018], their extension to flow-kick systems remains an open challenge, as it requires careful modifications to handle the interplay between flows and kicks.
This paper addresses these challenges by introducing invasion growth rates for a general class of flow-kick systems that include most forms of autonomous impulsive equations. Using these invasion growth rates, I provide two complementary approaches for establishing permanence: one based on Morse decompositions of the extinction set (collections of invariant sets that capture the fundamental dynamical behavior) and another using invasion graphs that characterize potential transitions between communities. Furthermore, these invasion growth rates are coupled with earlier work on identifying attractors in invariant hyperplanes [Ashwin et al., 1994, Rand et al., 1994] to identify extinction attractors. To illustrate how to use these results and provide new biological insights into how flow dynamics influences persistence, I analyze models for Lotka-Volterra systems experiencing periodic disturbances, models of serial passage experiments, and spatially structured consumer-resource dynamics with pulsed consumer reproduction. These applications demonstrate specific mechanisms through which the interplay between continuous flows and discrete kicks creates novel persistence outcomes that cannot emerge in purely continuous or purely discrete systems, including kick-induced coexistence in serial transfer experiments and extinction when reproductive events occur too frequently.
The framework developed in this paper provides a foundation for understanding how the complex interplay between continuous and discrete dynamics affects species persistence and community composition in natural systems. By extending invasion growth rate theory to flow-kick models, we gain new insights into predicting when species will coexist or go extinct in environments characterized by both gradual changes and sudden disturbances.
2. Model Framework
The models developed in this paper track the dynamics of $k$ species (or populations) through their densities $x=(x_{1},x_{2},\dots,x_{k})$ and $\ell$ auxiliary variables $y=(y_{1},y_{2},\dots,y_{\ell})$. This flexible structure allows us to represent various ecological complexities including species interactions, population structure (e.g. spatial, age, or genotypic distributions) and auxiliary processes (e.g. plant-soil feedbacks, seasonal forcing). The species densities take values in the nonnegative cone $[0,\infty)^{k}$ of the $k$-dimensional Euclidean space $\mathbb{R}^{k}$, while the auxiliary variables take values in the $\ell$-dimensional Euclidean space $\mathbb{R}^{\ell}$. Let $z=(x,y)$ correspond to the complete state of the system.
Three components govern the system’s impulsive dynamics: a vector field that determines the continuous-time dynamics (“the flow”), a mapping $z\to H(z)$ that governs discrete-time impulses (“the kick”), and a timing function $\tau$ that schedules when discrete events occur. The continuous dynamics are determined by the per capita growth rates $f_{i}(z)$ of each population $i=1,2,\dots,k$ and the rate of change $g_{i}(z)$ of the auxiliary variables for $i=1,2,\dots,\ell$. Thus, the equations of motion for the flow are:
(1) $\frac{dx_{i}}{dt}=$ $x_{i}f_{i}(z)\quad\mbox{ for }i\in[k]:=\{1,2,\dots,k\}$
$\frac{dy}{dt}=$ $g(z)\quad\mbox{where }z=(x,y).$
To ensure that the flow defined by (1) is uniquely determined, I make the following assumption:
A1:
There is an open set $\mathcal{K}\subset[0,\infty)^{k}\times\mathbb{R}^{\ell}$ in which the functions $f_{i},g_{j}:\mathcal{K}\to\mathbb{R}$ are locally Lipschitz for $i=1,\dots,k$, $j=1,\dots,\ell$. Consequently, there exists a unique solution $z.t$ to (1) for any initial condition $z=(x,y)\in\mathcal{K}$.
[…]
The timing of the kicks is allowed to be state dependent and is given by a continuous positive function $\tau$. To ensure that the model is well defined, I assume that solutions starting with $z$ are defined until at least time $\tau(z)$:
A2:
There is a closed set $\mathcal{S}\subset\mathcal{K}$ and a continuous positive function $\tau:\mathcal{S}\to(0,\infty)$ such that $z.t$ exists and lies in $\mathcal{K}$ for all $z\in\mathcal{S}$ and $0\leq t\leq\tau(z)$.
The set $\mathcal{S}$ corresponds to the possible states of the system immediately after a kick. After flowing for $\tau(z)$ units of time, the system enters the set
(2) $\mathcal{T}=\{z.\tau(z):z\in\mathcal{S}\},$
which contains all possible states just before a kick occurs. The system then experiences a kick determined by the mapping:
(3) $H:\mathcal{T}\to\mathcal{S}\mbox{ where }H(z)=(x_{1}F_{1}(z),\dots,x_{k}F_{k}(z),G(z))$
where $F_{i}(z)$ is the multiplicative change in the density of species $i$ and $G(z)$ updates the auxiliary variables. Concerning the kick mapping $H$, the following assumptions are made:
A3:
The functions $F_{i}:\mathcal{T}\to(0,\infty)$ for $i=1,\dots,k$ are continuous and positive. This ensures that kicks cannot cause immediate extinction of a species that is present.
A4:
The mapping $G:\mathcal{T}\to\mathbb{R}^{\ell}$ is continuous, ensuring continuous updates to the auxiliary variables.
To construct the full impulsive dynamics, I first consider the state changes between successive kicks given by iterating:
[…]
Given an initial condition $z\in\mathcal{S}$, the forward orbit of $z$ is the sequence
$\mathcal{O}^{+}(z)=\{z,\kappa^{1}(z),\kappa^{2}(z),\kappa^{3}(z),\dots\}$
where $\kappa^{n}(z)=(\underbrace{\kappa\circ\kappa\circ\dots\circ\kappa}_{n\mbox{ times}})(z)$ denotes $n$-fold composition of $\kappa$ with itself.
To extend the kick-to-kick dynamics to continuous time, I define the mapping $\Phi:[0,\infty)\times S\to\mathcal{K}$ as follows:
[…]
where $\mbox{where }T(0,z)$ $=0\mbox{ and }T(n,z)=T(n-1,z)+\tau(\kappa^{n-1}(z))\mbox{ for }n=1,2,3,\dots$
Here, $T(n,z)$ represents the cumulative time until the $n$-th kick occurs, which is constructed by summing the individual kick intervals $\tau$ for each state in the sequence. This formulation allows us to determine the state of the system at any continuous time $t$, not just at discrete kick times.
From this general flow-kick framework, we can recover two familiar dynamical systems as special cases. Pure ordinary differential equations can be recovered by choosing $\tau(z)=1$ for all $z$ and making $H$ the identity mapping. Discrete-time difference equations can be recovered by choosing the trivial vector field ($f_{i}(z)=0=g_{j}(z)$) and $\tau(z)=1$ for all $z$.
To illustrate the diversity of dynamics these models can capture, I present three ecological examples. The first example demonstrates the simplest type of flow-kick system where kicks reduce all populations by a fixed proportion. The second example simultaneously illustrates the use of state-dependent kick times $\tau$ and auxiliary variables $y$ to periodically switch flows. The final example shows how to use auxiliary variables to account for the population structure.
Figure 1. The flow-kick dynamics of: (A) Lotka-Volterra dynamics with periodic disturbances, (B) competing species experiencing serial passages, and (C) a two-patch model of resource-consumer interactions where the consumer reproduces at discrete time intervals. Parameters:
For model (A): $A={\tiny-\begin{pmatrix}1.0000000&0.7317073&0.2258065&0.1875000&0.3214286\\ 0.8571429&1.0000000&1.1612903&1.0250000&0.0892857\\ 0.6428571&0.8780488&1.0000000&0.9000000&0.2321429\\ 0.2142857&0.3658537&-11.6129032&1.0000000&1.0178571\\ 0.0000000&0.2682927&1.0967742&-0.9750000&1.0000000\end{pmatrix}}$, $r={\tiny\begin{pmatrix}1\\ 1\\ 1\\ 1\\ 1\end{pmatrix}}$, $\tau=24$, and $d=(0,0,0,0,0.9)$. For model (B): $\rho(1)=(0.4,0.1),\rho(2)=(0.02,0.1),\delta=0.1$ and $z(0)=(0.05,0.05,1)$. For model (C): $\alpha=(10,20)$, $\beta=(0.1,0.1)$,$a=(0.1,0.1)$, $m=(0.1,0.1)$,$\theta=0.5$,$\tau=52$,and $d=0.05$
2.1. Lotka-Volterra systems with periodic disturbances
To analyze how disturbances affect species coexistence, Huston [1979] studied Lotka-Volterra models with periodic density-independent population reductions. For a community of $k$ species, let $x_{i}$ denote the density of the $i$-th species, $r_{i}$ its intrinsic growth rate, and $a_{ij}$ the per-capita effect of species $j$ on the growth rate of species $i$. For this model, $\mathcal{C}=\mathcal{S}=\mathcal{T}=[0,\infty)^{k}$, meaning the state space for the continuous dynamics and the pre- and post-kick states are all the same. The population dynamics between disturbances follows the classic Lotka-Volterra equations:
$\frac{dx_{i}}{dt}=x_{i}\left(r_{i}+\sum_{j=1}^{k}a_{ij}x_{j}\right)$
Disturbances occur at regular intervals of length $\tau$, reducing the density of each species $i$ by a fixed proportion $d_{i}$. These instantaneous reductions can be expressed as the kick map $H:[0,\infty)^{k}\to[0,\infty)^{k}$ given by:
$H(x_{1},\dots,x_{k})=((1-d_{1})x_{1},\dots,(1-d_{k})x_{k})$
A sample simulation of this impulsive dynamical system for an empirically parameterized Lotka-Volterra model is shown in Figure 1A. The parameterization is based on four vascular plant species (Hypericum ellipticum, Juncus pelocarpus, Lysimachia terrestris, Dulichium arundinaceum) on the shoreline of Axe Lake, Ontario [Wilson and Keddy, 1986, Keddy and Shipley, 1989]. Wilson and Keddy [1986] estimated the interaction coefficients through pairwise competition experiments between all possible combinations of species. These experiments were initiated with two ramets per species and measured their relative biomass yields after three months of competition.
2.2. Competition models for serial passage experiments
Yi and Dean [2013] conducted competition experiments between two populations of E. coli: one chloramphenicol resistant (ClmR) and one resistant to tetracycline (TetR). Fluctuating selection was imposed by alternating sublethal doses of each antibiotic in a minimal glucose medium. When the cultures reached a certain optical density, they were diluted $100$ times in fresh medium with the alternate antibiotic. To model the dynamics of these competitors, Yi and Dean [2013] used a flow-kick model in which populations grew exponentially until reaching a critical density and then were diluted and introduced to another environment. In this model, $x=(x_{1},x_{2})$ corresponds to the densities of the two competitors. An auxiliary variable $y\in\{0,1\}$ keeps track of the environments, with chloramphenicol present in the environment 00 and tetracycline present in environment $1$. Let $\rho_{i}(y)$ be the per-capita growth rate of genotype $i$ in environment $y$. The flow dynamics represent uncoupled exponential growth of the two populations:
$\frac{dx_{i}}{dt}=$ $\rho_{i}(y)x_{i}\mbox{ for }i=1,2$
$\frac{dy}{dt}=$ 00 0
Without loss of generality, we set the critical density for a serial transfer to $1$. Let $\delta<1$ be the factor by which the population is diluted during a serial transfer. Then the post-kick state space $\mathcal{S}$ and the pre-kick state space $\mathcal{T}$ are:
$\mathcal{S}=\{x\in[0,\infty)^{2}:\sum_{i}x_{i}=\delta\}\times\{0,1\}\mbox{ and}\mathcal{T}=\{x\in[0,\infty)^{2}:\sum_{i}x_{i}=1\}\times\{0,1\}.$
The time to kick $\tau:\mathcal{S}\to(0,\infty)$ is implicitly defined as the unique solution to
$\sum_{i}x_{i}e^{\rho_{i}(y)\tau(z)}=1\mbox{ for }z=(x,y)\in\mathcal{S}.$
By the implicit function theorem, $\tau$ is a continuous function. The kick function $H:\mathcal{T}\to\mathcal{S}$ is:
$H(z)=(\delta x_{1},\delta x_{2},(y+1)\mbox{ mod }2)\mbox{ where }z=(x_{1},x_{2},y)$
This kick function dilutes both populations by the factor $\delta$ and switches the environment between states 00 and $1$. A sample simulation of this model is shown in Figure 1B.
2.3. Two Patch Pulsed Consumer-Resource Models
Pachepsky et al. [2008] introduced a flow-kick model of the dynamics of consumers and resources. In their model, the resource grows continuously and is continuously consumed throughout the year. Consumers accumulate consumed resources and die at a constant rate. Consumer reproduction occurs only once per year as a discrete pulse event proportional to the amount of resources consumed. These models capture the dynamics of consumers, such as barnacles, oysters, or clams, that feed continuously on plankton while breeding once per year.
Here, I consider a two-patch variant of this model to illustrate how auxiliary variables can capture spatial structure and biological feedback. Let $R_{i}$ denote the density of resources in patch $i=1,2$, $C_{i}$ the density of the consumer, and $B_{i}$ the amount of biomass accumulated per consumer. Unlike Pachepsky et al. [2008]’s model with logistic growth, resources grow at a constant rate $\alpha_{i}$ and have a per-capita loss rate $\beta_{i}$ in patch $i$. This chemostat-type dynamic can represent resources that aren’t only locally reproducing but also arriving from regions outside the focal patches, such as plankton arrival near two oyster reef patches. The consumer is sedentary and consumes the resource with a per-capita attack rate $a_{i}$ in patch $i$. It experiences a per-capita mortality rate $m_{i}$ in patch $i$. Between reproductive events, the dynamics follow:
(6) $\frac{dR_{i}}{dt}=$ $\alpha_{i}-\beta_{i}R_{i}-a_{i}R_{i}C_{i}$
$\frac{dC_{i}}{dt}=$ $-m_{i}C_{i}$
$\frac{dB_{i}}{dt}=$ $a_{i}R_{i}$
At fixed intervals of length $\tau>0$, consumers reproduce in proportion to their accumulated biomass, with offspring dispersing between patches with probability $d$. The instantaneous update of population densities follows:
(7) $(R_{1},R_{2},C_{1},C_{2},B_{1},B_{2})\mapsto(R_{1},R_{2},F_{1},F_{2},0,0)$
where $F_{1}=(1-d)C_{1}(1+\theta B_{1})+dC_{2}(1+\theta B_{2})$
and $F_{2}=dC_{1}(1+\theta B_{1})+(1-d)C_{2}(1+\theta B_{2})$
Here, $\theta$ represents the conversion efficiency of consumed resource to consumer individuals.
To facilitate analysis of this spatially structured system and align it with our general framework, I will express the model using a coordinate transformation that separates the total consumer density from its spatial distribution. This transformation will allow us to directly track the size of the consumer population (as our primary state variable), while treating the spatial structure and resource dynamics as auxiliary variables. To express this model in the coordinate system $z=(x,y)$, I define $x=C_{1}+C_{2}$ as the total density of consumers, $y_{1}=R_{1}$, $y_{2}=R_{2}$, $y_{3}=B_{1}$, $y_{4}=B_{2}$, and $y_{5}=C_{1}/x$ as the fraction of consumers in patch $1$. In these coordinates, the flow dynamics become:
(8) $\frac{dx}{dt}$ $=-m_{1}y_{5}x-m_{2}(1-y_{5})x$ $\frac{dy_{5}}{dt}=-m_{1}y_{5}+y_{5}(m_{1}y_{5}+m_{2}(1-y_{5}))$
$\frac{dy_{1}}{dt}$ $=\alpha_{1}-\beta_{1}y_{1}-a_{1}y_{1}xy_{5}$ $\frac{dy_{2}}{dt}=\alpha_{2}-\beta_{2}y_{2}-a_{2}y_{2}x(1-y_{5})$
$\frac{dy_{3}}{dt}$ $=a_{1}y_{1}$ $\frac{dy_{4}}{dt}=a_{2}y_{2}$
and the kick update map is:
(9) $H(z)=\left(\phi_{1}(z)+\phi_{2}(z),y_{1},y_{2},0,0,\phi_{1}(z)/(\phi_{1}(z)+\phi_{2}(z))\right).$
Here, $\phi_{1}(z)$ and $\phi_{2}(z)$ represent the transformed post-reproduction consumer densities in patches 1 and 2, respectively, corresponding to $F_{1}$ and $F_{2}$ in the original coordinates:
where $\phi_{1}(z)=(1-d)(1+\theta y_{3})xy_{5}+d(1+\theta y_{4})x(1-y_{5})$
and $\phi_{2}(z)=d(1+\theta y_{3})xy_{5}+(1-d)(1+\theta y_{4})x(1-y_{5}).$
The state space is $\mathcal{C}=\mathcal{S}=\mathcal{T}=[0,\infty)\times[0,\infty)^{4}\times[0,1]$ with $k=1$ (representing the species of the single consumer) and $\ell=5$ (representing the five auxiliary variables). A sample simulation of this model is shown in Figure 1C.
3. Extinction set and Invasion growth rates
For flow-kick models, the extinction set $\mathcal{S}_{0}:=\{z\in S|\prod_{i=1}^{n}x_{i}=0\}$ is the set of states where at least one species is extinct (i.e., it has a density equal to zero). The per-capita form of the flow and kick mapping imply that for any initial condition $z\in\mathcal{S}_{0}$, the future states $\kappa^{n}(z)$ remain in $\mathcal{S}_{0}$ for all time, capturing the principle of “no cats, no kittens” of closed ecological systems. The persistent set $\mathcal{S}_{+}=\mathcal{S}\setminus\mathcal{S}_{0}$ corresponds to states in which all species have positive densities. Assumptions A1 (via existence and uniqueness) and A3 (via positivity) imply that for any initial condition $z\in\mathcal{S}_{+}$, future states $\kappa^{n}(z)$ remain in $\mathcal{S}_{+}$ for all time. Notably, as with nearly all deterministic models, extinction can occur only asymptotically, not in finite time.
Permanence, defined more precisely in Section 4, corresponds to the extinction set $\mathcal{S}_{0}$ being a repellor, that is, if the system starts with all species present ($z\in\mathcal{S}_{+}$), then eventually it becomes bounded away from the extinction set. To understand permanence for flow-kick models, I use invasion growth rates, which measure the average per-capita growth rate of a population when introduced at infinitesimally small densities. The most mathematically convenient way to define these invasion growth rates is using ergodic theory. Section 3.1 provides some background on ergodic theory, introduces explicit formulas for invasion growth rates, highlights their key properties, and illustrates how to calculate these rates for the previously introduced examples.
3.1. Invasion growth rates
Biologically, invasion growth rates measure how a species would grow if introduced at very low densities into an established community. If the invasion growth rate is positive, the species can increase when rare and potentially establish itself. If negative, the species cannot invade. These rates are fundamental in predicting which species combinations can persist together and which cannot, essentially providing a mathematical foundation for understanding coexistence. To define invasion growth rates, we need some ergodic theory; see, for example,  [Mañé, 1983, Katok et al., 1995].
Consider a forward orbit $\{z,\kappa(z),\kappa^{2}(z),\dots\}$ of the kick-to-kick map $\kappa:\mathcal{S}\to\mathcal{S}$ (see (4)) for some initial condition $z\in\mathcal{S}$. For any continuous function $h:\mathcal{S}\to\mathbb{R}$ (an “observable”), it is natural to ask when the long-term average $\lim_{n\to\infty}\frac{1}{n}\sum_{m=0}^{n-1}h(\kappa^{m}(z))$ is well defined. Ergodic theory provides an answer in terms of invariant measures. A Borel probability measure $\mu$ on $\mathcal{S}$ is invariant for the kick-to-kick dynamics (4) if $\int h(z)\mu(dz)=\int h(\kappa(z))\mu(dz)$ for any continuous bounded function $h:\mathcal{S}\to\mathbb{R}$. That is, the average value of $h$ when randomly choosing the initial condition $z$ according to $\mu$ equals the average value of $h$ in the next time step.
An invariant probability measure $\mu$ is ergodic if it cannot be written as a non-trivial convex combination of two invariant probability measures, that is, if $\mu=\alpha\mu_{1}+(1-\alpha)\mu_{2}$ for two distinct invariant measures $\mu_{1},\mu_{2}$, then $\alpha=1$ or $\alpha=0$. Intuitively, the ergodic measures represent the most basic statistical behaviors of the system that cannot be further decomposed. The simplest example of an ergodic probability measure is a Dirac measure $\mu=\delta_{z^{*}}$ associated with an equilibrium $z^{*}\in\mathcal{S}$ of $\kappa$. This measure is defined by $\int h(z)\mu(dz)=h(z^{*})$ for every continuous bounded function $h:\mathcal{S}\to\mathbb{R}$. Similarly, if $z^{*}$ is a periodic point of period $N$ (i.e., $\kappa^{N}(z^{*})=z^{*}$), then the probability measure $\mu=\frac{1}{N}\sum_{n=0}^{N-1}\delta_{\kappa^{n}(z^{*})}$ defined by averaging along this periodic orbit is an ergodic measure. More generally, the ergodic theorem [Mañé, 1983, Chapter 6] implies that for every ergodic measure $\mu$ there exists an initial condition $\tilde{z}$ such that $\mu$ is determined by averaging along the forward orbit of $\tilde{z}$, that is, $\int h(z)\mu(dz)=\lim_{N\to\infty}\frac{1}{N}\sum_{n=0}^{N-1}h(\kappa^{n}(\tilde{z}))$ for all continuous bounded $h:\mathcal{S}\to\mathbb{R}$. In fact, such points $\tilde{z}$ make up almost all the support of $\mu$ (i.e., the set of points that do not satisfy this property has measure zero with respect to $\mu$). This means that for typical initial conditions within the support of an ergodic measure, time averages along trajectories converge to the average with respect to the measure.
Ergodic measures correspond to the long-term statistical behavior of subsets of species, what we might call ecological communities. To see why, for any subset of species $I\subset[k]$, let
$\mathcal{S}_{I}=\{z\in\mathcal{S}:x_{i}>0\mbox{ if and only if }i\in I\}$
be the set of states for which only the species in $I$ are present (i.e., have non-zero density). Since these sets $\mathcal{S}_{I}$ are invariant for kick-to-kick dynamics, an ergodic measure $\mu$ is uniquely associated with a set of species $I(\mu)\subseteq[k]$, which is the unique subset such that $\mu\left(\mathcal{S}_{I(\mu)}\right)=1$. I call $I(\mu)$ the species support of the ergodic measure $\mu$.
To define the average per-capita growth rate of species $i$, let $\pi_{i}:\mathcal{C}\to[0,\infty)$ be the projection of a state $z$ to the density of the $i$th species, i.e. $\pi_{i}(z)=x_{i}$ for $z=(x,y)\in\mathcal{C}$. For $z$ with $x_{i}>0$, the fundamental theorem of calculus implies that $\ln\frac{\pi_{i}(\kappa(z))}{\pi_{i}(z)}$ equals
[…]
This represents the per-capita growth rate of species $i$ over one complete flow-kick cycle starting in state $z$. Notice that $r_{i}(z)$ is well defined even if $x_{i}=0$. In this case, $r_{i}(z)$ describes the rate at which species $i$ would increase if introduced at infinitesimally small densities.
For any probability measure $\mu$, we define
(11) $r_{i}(\mu)=\int r_{i}(z)\mu(dz)\mbox{, the average per-capita growth rate of species $i$ with respect to $\mu$.}$
$r_{i}(\mu)$ measures on average how species $i$ would grow if, in the long-term, the state of system follows the distribution given by $\mu$. If $\mu$ is an ergodic measure, Birkhoff’s ergodic theorem implies that
(12) $\lim_{n\to\infty}\frac{1}{n}\sum_{m=0}^{n-1}r_{i}(\kappa^{m}(z))=r_{i}(\mu)$
[…]
In the case of an ergodic measure $\mu$, we intuitively expect the average per-capita growth rate of a species supported by $\mu$ to be zero. This is because species densities that remain bounded away from both zero and infinity must have a long-term average growth rate of zero. The following lemma confirms this intuition.
Lemma 3.1.
Let $\mu$ be an ergodic probability measure for the kick-to-kick map $\kappa$ with compact support. Then $r_{i}(\mu)=0$ for all $i\in I(\mu)$.
The proof strategy follows from an argument given for ODE models without auxiliary variables $y$ found in [Schreiber, 2000, Lemma 5.1].
Proof.
Since $\mu(\mathcal{C}_{I(\mu)})=1$, Birkhoff’s Ergodic Theorem implies that there exists an invariant Borel set $U\subseteq\mathcal{C}_{I(\mu)}$ such that $\mu(U)=1$ and
(13) $\lim_{n\to\infty}\frac{1}{n}\sum_{m=0}^{n-1}r_{i}(\kappa^{m}(z))=r_{i}(\mu)$
whenever $z\in U$. Choose an open set $V$ such that its closure $\overline{V}$ is contained in $\mathcal{C}_{I(\mu)}$, $\overline{V}$ is compact, and $\mu(V\cap U)>0$. By the Poincaré recurrence theorem, there exists $z\in V\cap U$ and an increasing sequence of integers $n_{m}\uparrow\infty$ such that $\kappa^{n_{m}}(z)\in V$ for all $m\geq 1$. Since $\overline{V}$ is compact, there exists a $\delta>0$ such that
(14) $1/\delta\leq\pi_{i}(\kappa^{n_{m}}(z))\leq\delta$
for all $m$. As
$\frac{1}{n}\ln\frac{\pi_{i}(\kappa^{n}(z))}{\pi_{i}(z)}=$ $\frac{1}{n}\ln\prod_{m=0}^{n-1}\frac{\pi_{i}(\kappa^{m+1}(z))}{\pi_{i}(\kappa^{m}(z))}=\frac{1}{n}\sum_{m=0}^{n-1}\ln\frac{\pi_{i}(\kappa^{m+1}(z))}{\pi_{i}(\kappa^{m}(z))}$
$=$ $\frac{1}{n}\sum_{m=0}^{n-1}\left(\int_{0}^{\tau(\kappa^{m}(z))}f_{i}(\kappa^{m}(z).t)dt+\ln F_{i}(\kappa^{m}(z).\tau(\kappa^{m}(z))\right)=\frac{1}{n}\sum_{m=0}^{n-1}r_{i}(\kappa^{m}(z))$
Since $\pi_{i}(\kappa^{n_{m}}(z))$ is bounded both above and below by positive constants as shown in (14), $\frac{\pi_{i}(\kappa^{n_{m}}(z))}{\pi_{i}(z)}$ must approach zero as $m\to\infty$. Combined with equation (13), this proves that
$r_{i}(\mu)=\lim_{m\to\infty}\frac{1}{n_{m}}\sum_{s=0}^{n_{m}-1}r_{i}(\kappa^{s}(z))=0.$
for all species $i\in I(\mu)$. ∎
Let us revisit the three examples introduced in Section 2 and derive their invasion growth rates for the ergodic measures supported on the extinction set.
3.2. $r_{i}(\mu)$ for Lotka-Volterra systems with periodic disturbances
Consider the disturbed Lotka-Volterra models introduced in Section 2.1. For these systems, given an ergodic measure $\mu(dx)$, the average per-capita growth rate of species $i$ is:
$r_{i}(\mu)=\int\left(\int_{0}^{\tau}f_{i}(x.t)dt\right)\mu(dx)+\ln(1-d_{i}),$
where $f_{i}(x)=\sum_{j}a_{ij}x_{j}+r_{i}$ represents the per-capita growth rate of species $i$.
These average per-capita growth rates typically admit explicit solutions. To see why, let $x_{i}^{*}(\mu)$ denote the average density of species $i$ with respect to $\mu$:
$x_{i}^{*}(\mu)=\frac{1}{\tau}\int\left(\int_{0}^{\tau}\pi_{i}(z.t)dt\right)\mu(dx)$
where $\pi_{i}:\mathcal{S}\to\mathbb{R}$ is the projection onto the density of species $i$, that is, $\pi_{i}(z)=x_{i}$ for $z=(x,y)$. Ecologically, $x_{i}^{*}(\mu)$ represents the time-averaged density of species $i$ in the community represented by $\mu$. These average densities allow us to calculate invasion growth rates without needing to track the detailed temporal dynamics between disturbances. Taking advantage of the linearity of the per-capita growth rate, $r_{i}(\mu)$ can be expressed in terms of these average densities:
(15) $r_{i}(\mu)=\tau\left(\sum_{j}a_{ij}x_{j}^{*}(\mu)+r_{i}\right)+\ln(1-d_{i}).$
The average densities $x_{i}^{*}(\mu)$ can be determined by observing that: (i) Lemma 3.1 implies that $r_{i}(\mu)=0$ for all species $i\in I(\mu)$ supported by $\mu$, and (ii) $x_{i}^{*}(\mu)=0$ for all species $i\notin I(\mu)$ absent from the support of $\mu$. These properties yield a system of linear equations:
(16) $x_{i}^{*}(\mu)$ $=0$ for $\text{for }i\notin I(\mu)$
$\sum_{j}a_{ij}x_{j}^{*}(\mu)+r_{i}+\frac{\ln(1-d_{i})}{\tau}$ $=0$ for $\text{for }i\in I(\mu)$
When this system admits a unique solution (which occurs generically), explicit expressions for all average per-capita growth rates $r_{i}(\mu)$ can be obtained by substituting the solved values of $x_{i}^{*}(\mu)$ into equation (15).
3.3. $r_{i}(\mu)$ for Competition models for serial passage experiments
Consider the competition model in Section 2.2 for serial passage experiments. The kick-to-kick function for this model is:
$\kappa(z)=(x_{1}e^{\rho_{1}(y)\tau(z)},x_{2}e^{\rho_{2}(y)\tau(z)},(y+1)\mbox{mod }2)\mbox{ where }z=(x_{1},x_{2},y)$
and $\tau$ is implicitly defined by:
$\sum_{i}x_{i}e^{\rho_{i}(y)\tau(z)}=\frac{1}{\delta}.$
The extinction set $\mathcal{S}_{0}$ for the kick-to-kick map $\kappa$ consists of four points $(x_{1},x_{2},y)=(\delta,0,0),(0,\delta,0),(\delta,0,1),(0,\delta,1)$, which correspond to two period-two orbits: $(\delta,0,0)\rightarrow(\delta,0,1)\rightarrow(\delta,0,0)$ and $(0,\delta,0)\rightarrow(0,\delta,1)\rightarrow(0,\delta,0)$. These orbits represent scenarios where only one species persists, while the environmental variable $y$ alternates between 0 and 1. The only ergodic invariant measures in the extinction set are Dirac measures supported on these periodic orbits: $\mu_{1}(dz)$ with $\mu_{1}(\{(\delta,0,0)\})=\mu_{1}(\{(\delta,0,1)\})=1/2$ and $\mu_{2}(dz)$ with $\mu_{2}(\{(0,\delta,0)\})=\mu_{2}(\{(0,\delta,1)\})=1/2$.
For an ergodic measure $\mu$, the average per-capita growth rate of species $i$ is:
$r_{i}(\mu)=\int\rho_{i}(y)\tau(z)\mu(dz)+\ln\delta$
To evaluate $r_{i}(\mu_{1})$, note that $\tau(\delta,0,y)=\frac{1}{\rho_{1}(y)}\ln\frac{1}{\delta}$. This yields the following expressions for $r_{i}(\mu_{1})$ and, through analogous reasoning, $r_{i}(\mu_{2})$:
(17) ${}r_{i}(\mu_{1})=\frac{\ln\frac{1}{\delta}}{2}\sum_{y\in\{0,1\}}\left(\frac{\rho_{i}(y)}{\rho_{1}(y)}-1\right)\mbox{ and }r_{i}(\mu_{2})=\frac{\ln\frac{1}{\delta}}{2}\sum_{y\in\{0,1\}}\left(\frac{\rho_{i}(y)}{\rho_{2}(y)}-1\right)\mbox{ for }i=1,2.$
These expressions show how invasion growth rates depend on the ratio of growth rates in each environment, weighted by the time spent in each environment. A species can invade if, on average in both environments, its growth rate relative to the resident species exceeds the penalty for dilution.
3.4. $r_{i}(\mu)$ for the Two Patch Pulsed Consumer-Resource Models
Consider the two-patch model of consumer-resource interactions from Section 2.3. The state space for the kick-to-kick map is $\mathcal{S}=[0,\infty)\times[0,\infty)^{4}\times[0,1]$ and the extinction set is $\mathcal{S}_{0}=\{(0,y)\in\mathcal{S}\}$, which corresponds to the extinction of the consumer.
For an initial condition $z\in\mathcal{S}_{0}$ on the extinction set, the $y_{1},y_{2},y_{3},y_{4}$ coordinates of the forward orbit $\{z,\kappa(z),\kappa^{2}(z),\dots\}$ of the kick-to-kick map converge to $(\alpha_{1}/\beta_{1},\alpha_{2}/\beta_{2},0,0)$. At these equilibrated values, the kick-to-kick dynamics of the coordinate $y_{5}$ corresponds to how the fraction of consumers in patch $1$ changes in the limit of an infinitesimally small consumer population. These dynamics are determined by the fractional linear map:
$y_{5}\mapsto\frac{d\gamma_{1}y_{5}+(1-d)\gamma_{2}(1-y_{5})}{\gamma_{1}y_{5}+\gamma_{2}(1-y_{5})},\text{ where }\gamma_{i}=e^{-m_{i}\tau}(1+\theta\tau a_{i}\alpha_{i}/\beta_{i})$
that has the unique globally stable equilibrium:
(18) $y_{5}^{*}=\frac{-2\gamma_{2}+d\gamma_{2}+d\gamma_{1}+\sqrt{d^{2}\gamma_{2}^{2}-8d\gamma_{1}\gamma_{2}+2d^{2}\gamma_{1}\gamma_{2}+4\gamma_{1}\gamma_{2}+d^{2}\gamma_{1}^{2}}}{2\left(\gamma_{1}-\gamma_{2}\right)}.$
Therefore, the only ergodic measure $\mu$ for the kick-to-kick map supported in the extinction set $\mathcal{S}_{0}$ is the Dirac measure at equilibrium $z^{*}=(0,\alpha_{1}/\beta_{1},\alpha_{2}/\beta_{2},0,0,y_{5}^{*})$. The average per-capita growth rate of the consumers at this ergodic measure is:
(19) $r_{i}(\mu)=y_{5}^{*}\ln\gamma_{1}+(1-y_{5}^{*})\ln\gamma_{2}$
This average per-capita growth rate and the equilibrium value $y_{5}^{*}$ relate to the dominant eigenvalue and eigenvector of the extinction equilibrium with respect to the original $(C_{1},C_{2})$ coordinate system. To see why, the derivative matrix of the kick-to-kick consumer dynamics at the extinction equilibrium $(R_{1},R_{2},C_{1},C_{2})=(\alpha_{1}/\beta_{1},\alpha_{2}/\beta_{2},0,0)$ is:
(20) $\begin{pmatrix}(1-d)\gamma_{1}&d\gamma_{2}\\ d\gamma_{1}&(1-d)\gamma_{2}\end{pmatrix}$
The dominant eigenvalue of this matrix equals $e^{r_{i}(\mu)}$ and the associated eigenvector is $\begin{pmatrix}y_{5}^{*}&1-y_{5}^{*}\end{pmatrix}^{\intercal}$ where denotes the transpose. This connection between invasion growth rates and eigenvalues of the linearized system demonstrates how the flow-kick framework extends classical stability analysis. The dominant eigenvalue $e^{r_{i}(\mu)}$ determines whether small consumer populations grow ($r_{i}(\mu)>0$) or decline ($r_{i}(\mu)<0$), while the associated eigenvector gives the spatial distribution that would be observed during this growth or decline.
4. Main results
To characterize permanence using invasion growth rates, I take two complementary approaches. The more general approach uses Morse decompositions in the extinction set $\mathcal{S}_{0}$, relying on earlier topological characterizations of permanence [Butler et al., 1986, Hofbauer and So, 1989, Garay, 1989]. The second approach uses invasion graphs [Hofbauer and Schreiber, 2022] that identify possible trajectories that connect different subsets of species according to invasion growth rates. These invasion graphs provide a biologically meaningful way to construct Morse decompositions.
To state the main results, I first assume that the dynamics of the kick-to-kick map are uniformly bounded from above, meaning the system is dissipative. Let $\|z\|=\sum_{i}|x_{i}|+\sum_{j}|y_{j}|$ denote the $L^{1}$ norm for $z=(x,y)\in[0,\infty)^{k}\times\mathbb{R}^{\ell}$. This norm corresponds to the total density of all species when restricted to the coordinates $x$, that is, $\|x\|=\sum_{i}x_{i}$ for $x\in[0,\infty)^{k}$.
A5:
The kick-to-kick map $\kappa$ is dissipative: There exists $K>0$ such that
$\limsup_{n\to\infty}\|\kappa^{n}(z)\|\leq K$
for all $z\in\mathcal{S}$.
The dissipativeness of $\kappa$ implies that there is a global attractor $\Gamma\subset\mathcal{S}$ for $\kappa$. To define this precisely, I recall some definitions from dynamical systems. For a compact set $C\subset\mathcal{S}$, its $\omega$ limit set is defined as:
$\omega(C):=\cap_{N\geq 1}\overline{\cup_{n\geq N}\kappa^{n}(C)}$
where $\kappa^{n}(C)$ represents the $n$-th iterate of $\kappa$ applied to the set $C$ i.e. $\kappa(C)=\cup_{z\in K}\kappa^{n}(z)$. The $\omega$-limit set characterizes the long-term behavior of the forward orbits of the kick-to-kick dynamics. A compact set $\Gamma\subset\mathcal{S}$ is an attractor if it has a compact neighborhood $U$ such that all trajectories starting in $U$ eventually converge uniformly to $\Gamma$ i.e. $\omega(U)=\Gamma$. An attractor $\Gamma$ is a global attractor if $\omega(z)\subset\Gamma$ for all $z\in\mathcal{S}$.
Permanence ensures that no species goes extinct regardless of initial conditions, provided that all species are initially present. Under Assumption A5, the kick-to-kick map is permanent if there is a minimal density $M>0$ such that for any initial condition $z\in\mathcal{S}_{+}$ with all species present,
(21) $\liminf_{n\rightarrow\infty}\pi_{i}(\kappa^{n}(z))\geq M\mbox{ for all }i=1,2,\dots,k$
where $\pi_{i}(z)$ corresponds to the density of species $i$. Permanence implies that if all species are initially present, they will coexist in the long term, even when faced with rare but large perturbations or frequent small perturbations [Schreiber, 2006]. In terms of attractors, permanence corresponds to a positive attractor $A\subset\mathcal{S}_{+}$ whose basin of attraction is $\mathcal{S}_{+}$, which means $\omega(z)\subset A$ for all $z\in\mathcal{S}_{+}$.
The following lemma proves that permanence of the kick-to-kick map implies the permanence for the continuous-time flow-kick dynamics. A proof is given at the end of Section 7.
Lemma 4.1.
Assume assumptions A1A5 hold. If the kick-to-kick map $\kappa$ is permanent, then the flow-kick dynamics $\Phi$, as defined by (5), is dissipative and permanent: There are constants $\widetilde{K}\geq\widetilde{M}>0$ such that
$\limsup_{t\rightarrow\infty}\|\Phi(t,z)\|\leq\widetilde{K}$
whenever $z\in\mathcal{S}$ and
$\liminf_{t\rightarrow\infty}\pi_{i}(\Phi(t,z))\geq\widetilde{M}\mbox{ for all }i=1,2,\dots,k$
whenever $z\in\mathcal{S}_{+}$.
4.1. Permanence and Extinction via Morse decompositions
I first state conditions for permanence using Morse decompositions, which provide a way to decompose the dynamics on the extinction set into fundamental building blocks. To define a Morse decomposition, I need to introduce the concept of a negative orbit. A sequence $\mathcal{O}^{-}=\{z(0),z(-1),z(-2),\dots\}\subset\mathcal{S}$ is a negative orbit of $\kappa$ if $\kappa(z(-k))=z(-k+1)$ for all $k=1,2,\dots$. That is, the sequence is consistent with the kick-to-kick dynamics when moving backward in time. As the mapping $\kappa$ need not be a homeomorphism, a negative orbit may not exist for a given initial condition $z\in\mathcal{S}$ or may not be unique when it does exist. However, since the global attractor $\Gamma$ is invariant (i.e., $\kappa(\Gamma)=\Gamma$), negative orbits exist for all initial conditions in $\Gamma$ and can be chosen to lie entirely in $\Gamma$. The $\alpha$-limit set of a negative orbit $\mathcal{O}_{-}$ is defined as:
$\alpha(\mathcal{O}_{-})=\{\widetilde{z}\in\mathcal{S}:\lim_{n_{k}\to\infty}z(-n_{k})=\widetilde{z}\mbox{ for some sequence }n_{k}\uparrow\infty\}$
The $\alpha$-limit set characterizes the asymptotic behavior of the negative orbit in backward time.
Let $\Gamma_{0}=\Gamma\cap\mathcal{S}_{0}$ be the restriction of $\Gamma$ to the extinction set and $\Gamma_{+}=\Gamma\cap\mathcal{S}_{+}$ be the restriction of $\Gamma$ to where all species persist.
I will now define a Morse decomposition that decomposes the extinction set $\Gamma_{0}$ into a finite number of compact invariant sets with specific dynamical properties. A set $A\subset\mathcal{S}$ is invariant for $\kappa$ if $\kappa(A)=A$, and an invariant set $A$ is isolated if there exists a closed neighborhood $U$ of $A$ such that $A$ is the maximal invariant set in $U$.
Definition 4.2.
A collection of sets =$\mathcal{M}=\{M_{1},M_{2},...,M_{\ell}\}$ is a Morse decomposition for $\Gamma_{0}$ if $M_{1},M_{2},...,M_{\ell}$ are pairwise disjoint isolated invariant compact sets, called Morse sets, such that $M_{i}\subset\Gamma_{0}$ and for every $z\in\Gamma_{0}\backslash\cup_{m=1}^{\ell}M_{m}$ there are integers $i<j$ such that $\omega(z)\subset M_{j}$ and $\alpha(\mathcal{O}_{-})\subset M_{i}$ for all negative orbits $\mathcal{O}^{-}=\{z(n)\}_{n=-\infty}^{\infty}\subset\Gamma_{0}$ with $z(0)=z$.
Intuitively, a Morse decomposition divides the extinction set $\Gamma_{0}$ into isolated invariant pieces that are ordered in such a way that dynamics can only move from lower-indexed to higher-indexed pieces, creating a hierarchical structure of the flow. Morse decompositions always exist but are not necessarily unique. Trivially, one Morse decomposition of $\Gamma_{0}$ is $\{\Gamma_{0}\}$ itself. However, Morse decompositions become more useful when they are more refined than this trivial decomposition.
The first main theorem uses Morse decompositions to provide conditions in terms of invasion growth rates that ensure permanence.
Theorem 4.3 (Permanence via Morse decompositions).
Assume assumptions A1A5 hold. Let =$\mathcal{M}=\{M_{1},M_{2},\dots M_{\ell}\}$ be a Morse Decomposition for $\Gamma_{0}$, where $\Gamma_{0}=\mathcal{S}_{0}\cap\Gamma$ and $\Gamma$ is the global attractor for the kick-to-kick map $\kappa$. If for each Morse set $M_{m}$
(22) $\max_{i}r_{i}(\mu)>0\mbox{ for all invariant measures $\mu$ with }\mu(M_{m})=1,$
then (5) is permanent.
Biologically, the condition (22) means that for each invariant set $M_{m}$ in the extinction set, there is at least one missing species that can invade. This condition implies that each Morse component $M_{m}$ is repelling to trajectories with all species initially present. By the minimax theorem, condition (22) is equivalent to the following condition: there exist positive weights $p_{1},\dots,p_{k}$ (possibly depending on $m$) such that
(23) $\sum_{i=1}^{k}p_{i}r_{i}(\mu)>0\mbox{ for all \emph{ergodic} measures $\mu$ with $\mu(M_{m})=1$}$
This equivalent condition is sometimes known as the Hofbauer condition, as it was first introduced by Hofbauer [1981] for permanence of ODEs in the special case of the trivial Morse decomposition (i.e., $\ell=1$ and $M_{1}=\Gamma_{0}$). The equivalence of conditions (22) and (23) was first observed by Garay and Hofbauer [2003].
The proof of Theorem 4.3 is based on a topological characterization of permanence due to Hofbauer and So [1989] and the following lemma. The lemma implies that for any initial state with all species present, the average per-capita growth rates associated with its $\omega$-limit set are non-positive and the per-capita growth rates associated with its $\alpha$-limit sets are non-negative. For continuous-time models, this lemma provides a means of simplifying the proof of Theorem 4.3 for models without auxiliary variables [Schreiber, 2000] and providing a more direct approach to proving the invasion graph characterization of permanence [Hofbauer and Schreiber, 2022].
Lemma 4.4.
For any $z\in\mathcal{S}_{+}$, there exists an invariant probability measure $\mu$ such that $\mu(\omega(z))=1$ and $r_{i}(\mu)\leq 0$ for all $i\in[k]$. For any negative orbit $\mathcal{O}^{-}\subset\Lambda\cap\mathcal{S}_{+}$, there exists an invariant probability measure $\mu$ such that $\mu(\alpha(\mathcal{O}^{-}))=1$ and $r_{i}(\mu)\geq 0$ for all $i\in[k]$.
I now provide partial converses to Theorem 4.3, giving sufficient conditions for the existence of an extinction-bound attractor and an attractor on the extinction set. Both conditions require that the kick-to-kick map $\kappa$ has additional smoothness.
Proposition 4.5 (Extinction bound trajectories).
Assume the kick-to-kick map $\kappa$ is twice continuously differentiable, and assumptions A1A5 hold. If there exists an ergodic measure $\mu$ such that $\mu(\mathcal{S}_{0})=1$ and $r_{i}(\mu)<0$ for all $i\notin I(\mu)$, then there exists $z\in\mathcal{S}_{+}$ such that $\omega(z)\subset\mathcal{S}_{0}$. In particular, the system is not permanent.
This proposition states that if there is a community of species in the extinction set (represented by an ergodic measure $\mu$) that cannot be invaded by any of the missing species (that is, all missing species have negative invasion growth rates), then there exists at least one trajectory starting with all species present that leads to the extinction of some species. This trajectory serves as a counterexample to permanence. The second condition follows from [Ashwin et al., 1994, Theorem 2.12] and Assumption A3.
Proposition 4.6 (Extinction attractors).
Assume that the kick-to-kick map $\kappa$ is twice continuously differentiable, and assumptions A1A5 hold. If
  1. (1)
    there is a subset of species $I\subset[k]$ and an attractor $A\subset\mathcal{S}_{I}$ for the kick-to-kick $\kappa$ dynamics restricted to $\mathcal{S}_{I}$, and
  2. (2)
    $r_{i}(\mu)<0$ for all $i\notin I$ and for all ergodic measures $\mu$ with $\mu(A)=1$
then $A$ is an attractor for the kick-to-kick map $\kappa:\mathcal{S}\to\mathcal{S}$.
This proposition provides a systematic way to identify extinction attractors by: (1) verifying the permanence conditions of Theorem 4.3 for a subset $I$ of species, and (2) checking that no missing species can invade (the non-invasibility condition $\max_{i\notin I}r_{i}(\mu)<0$). When both conditions are met, the attractor $A$ persists in the full system.
4.2. Permanence and Extinction via Invasion Graphs
To help construct the Morse decompositions needed by Theorem 4.3, Hofbauer and Schreiber [2022] introduced invasion graphs as a method to identify Morse decompositions for ODE models. Invasion graphs characterize all possible transitions between communities due to single- or multiple-species invasions. Here, I extend the invasion graph approach to flow-kick systems and show how it can be used to characterize permanence. To define the invasion graph, I need the following additional assumptions:
A6:
For each ergodic invariant Borel probability measure $\mu$ supported by $\mathcal{S}_{0}$, $r_{j}(\mu)\neq 0$ for all $j\notin I(\mu)$, and
A7:
$\mathop{\hbox{sgn}}r_{j}(\mu)=\mathop{\hbox{sgn}}r_{j}(\nu)$ for any two ergodic measures $\mu,\nu$ with $I(\mu)=I(\nu)$ and all $j$.
Assumption A6 requires that invasion growth rates $r_{i}(\mu)$ be non-zero for species not supported by $\mu$. This non-degeneracy condition ensures that each species either can or cannot invade a given community. This assumption holds typically for dissipative Lotka-Volterra systems or systems with a finite number of ergodic measures. Assumption A7 ensures that all ergodic measures with the same species support agree on which missing species can invade, allowing us to define invasion growth rates for communities rather than for specific ergodic measures. This assumption automatically holds when each face supports at most one invariant probability measure (e.g., there is a unique equilibrium, periodic orbit, or quasi-periodic motion in a given face). Due to their time-averaging property, assumption A7 also holds for the periodically disturbed Lotka-Volterra systems introduced in Section 2.1.
However, A7 can fail in some important situations. For example, McGehee and Armstrong [1977] demonstrated coexistence in the sense of a positive attractor but not permanence for a three-species system consisting of two predators and one prey species. In their example, a predator-prey pair coexists along a stable limit cycle at which the other predator can invade ($r_{i}(\mu)>0$ for the ergodic measure $\mu$ supported by the limit cycle), but the predator-prey pair also supports an unstable equilibrium at which the other predator cannot invade ($r_{i}(\nu)<0$ for the Dirac measure $\nu$ supported by this equilibrium).
Given assumption A7, I can uniquely define:
$r_{i}(I)=\mathop{\hbox{sgn}}r_{i}(\mu)\mbox{ whenever }I(\mu)=I.$
Let $\mathcal{C}$ be the set of all communities: all subsets $I$ of $[k]$ such that $I=I(\mu)$ for some ergodic measure $\mu$. Let $|\mathcal{C}|$ be the number of elements in $\mathcal{C}$. Note that $\mathcal{C}$ may not include all possible subsets of $[k]$, but only those that can actually be achieved as the support of some ergodic measure. Following Hofbauer and Schreiber [2022], the invasion scheme $\mathbb{IS}$ is the table of signs of invasion growth rates $\{(r_{i}(I))_{i\in[k]}:I\in\mathcal{C}\}$. This scheme can be represented as a matrix $|\mathcal{C}|\times k$ where the rows correspond to communities and the columns correspond to species, with entries that indicate whether each species can invade ($+1$), cannot invade ($-1$), or is already present (00) in each community.
Invasion growth rates are useful for identifying possible limit sets $\alpha$ and $\omega$ for an orbit of kick-to-kick dynamics. The following crucial lemma is proved in Section 8. Roughly, the lemma states that if a trajectory starting with a set of species converges to a set supporting only a subset of these species, then the missing species must have negative invasion growth rates in this limit. Conversely, if a negative orbit starting with a set of species converges in backward time to a subset of these species, then the missing species must have positive invasion growth rates in the limit into the deep past. The proof is based on Lemma 4.4 and is given in Section 8.
Lemma 4.7.
Assume A1A7 hold. If $z$ is an initial condition in $\mathcal{S}_{I}$ for some $I\subset[k]$ and $\omega(z)$ for the kick-to-kick map $\kappa$ lies in $\mathcal{S}_{J}$ for some $J\subset I$, then $J\in\mathcal{C}$ and $r_{j}(J)<0$ for all $j\in I\setminus J$. Alternatively, if $\mathcal{O}^{-}=\{z(n)\}_{n=-\infty}^{0}\subset\Gamma$ is a negative orbit where $z(0)\in\mathcal{S}_{I}$ for some $I\subset[k]$ and $\alpha(\mathcal{O}^{-})\subset\mathcal{S}_{J}$ for some $J\subset I$, then $J\in\mathcal{C}$ and $r_{j}(J)>0$ for all $j\in I\setminus J$.
Based on this lemma, signs of invasion growth rates can identify possible transitions between communities. The invasion graph $\mathbb{IG}$ is the directed graph whose vertices are the communities $\mathcal{C}$ and for which there is a directed edge from community $I$ to community $J$ if:
  • $I\neq J$ (communities are distinct),
  • $r_{j}(I)>0$ for all $j\in J\setminus I$ (all species in $J$ but not in $I$ can invade community $I$), and
  • $r_{i}(J)<0$ for all $i\in I\setminus J$ (all species in $I$ but not in $J$ cannot reinvade community $J$).
This directed edge represents a possible transition from community $I$ to community $J$ through ecological invasion processes.
Importantly, Lemma 4.7 implies that if a trajectory has its $\alpha$-limit set in $\mathcal{S}_{I}$ for a proper subset $I\subset[k]$ and its $\omega$-limit set in $\mathcal{S}_{J}$ for another proper set $J\neq I$, then there is a directed edge from $I$ to $J$ on the invasion graph. The opposite need not be true: a directed edge from $I$ to $J$ in the invasion graph does not guarantee the existence of an actual trajectory whose $\omega$-limit set lies in $\mathcal{S}_{J}$ and whose $\alpha$-limit set lies in $\mathcal{S}_{I}$.
When the invasion graph is acyclic (i.e., contains no directed cycles), the following theorem characterizes permanence. It requires that every community $I\in\mathcal{C}$ with some missing species has at least one missing species with a positive invasion growth rate, that is, $\max_{i\notin I}r_{i}(I)>0$. This theorem generalizes Hofbauer and Schreiber [2022, Theorem 1], which corresponds to the special case of ordinary differential equations.
Theorem 4.8.
Assume that A1A7 hold and that $\mathbb{IG}$ is acyclic. If $\max_{i\notin I}r_{i}(I)>0$ for each $I\in\mathcal{C}\setminus\{[k]\}$, then (5) is permanent.
In the development of the theory of coexistence based on invasion growth rates for competitive communities, Chesson [1994] introduced the concept of $-i$ communities: communities that arise when only species $i$ is missing. More precisely, a community $I\subset[k]$ is called an $-i$ community if: (1) $i\notin I$, and (2) $r_{j}(I)<0$ for all $j\in[k]\setminus(I\cup\{i\})$. Biologically, an $-i$ community is a subcommunity that cannot be invaded by any species except possibly species $i$. By Lemma 4.7, if $z=(x,y)\in\mathcal{S}$ is an initial condition with only species $i$ absent (i.e., $x_{i}=0$ and $x_{j}>0$ for $j\neq i$) and $\omega(z)\subset\mathcal{S}_{I}$ for some $I\subset[k]$, then $I$ must be an $-i$ community. Theorem 4.8 implies the following corollary:
Corollary 4.9.
Assume A1A7 hold and $\mathbb{IG}$ is acyclic. If $r_{i}(I)>0$ for all $-i$ communities with $i\in[k]$, then (5) is permanent.
This corollary provides a way to check permanence by focusing only on $-i$ communities, which are often easier to identify than the full invasion graph.
When the invasion graph is acyclic, Proposition 4.5 implies that the permanence condition in Theorem 4.8 is sharp. That is, if there is a community $I$ with some missing species and all the missing species have negative per-capita growth rates (i.e., $\max_{i\not\in I}r_{i}(I)<0$), then the system has an extinction-bound trajectory and is not permanent. Invasion graphs also provide a way to identify extinction attractors:
Corollary 4.10 (Extinction attractors via Invasion Graphs).
Assume A1A7 hold and the kick-to-kick map is twice continuously differentiable. Let $I\in\mathcal{C}$ be such that:
  1. (1)
    The invasion graph restricted to $I$ is acyclic and satisfies $\max_{j\in I\setminus J}r_{j}(J)>0$ for any $J\subsetneq I$ with $J\in\mathcal{C}$, and
  2. (2)
    $r_{i}(I)<0$ for all $i\notin I$,
then there is an attractor $A$ for the kick-to-kick map such that $A\subset\mathcal{S}_{I}$ and there exists a positive continuous function $\eta:\mathcal{S}_{I}\to(0,\infty)$ such that $\omega(z)\subset A$ whenever $\pi_{i}(z)>0$ for $i\in I$ and $\pi_{i}(z)<\eta(z)$ for $i\notin I$.
The last statement has a clear ecological interpretation: whenever all species in the community $I$ are initially present, any sufficiently small introduction of species outside $I$ will fail to establish, and the system will return to the attractor $A$ where only species in $I$ persist. This provides a mathematical formalization of what ecologists call an “invasion-resistant” community.
5. Applications
To illustrate how to apply the theoretical results from Section 4, I revisit the three examples introduced in Section 2 to identify the conditions for extinction and permanence. Each example demonstrates different aspects of the theory while raising open research questions and providing new biological insights.
Figure 2. The impact of disturbances on an empirically parameterized Lotka-Volterra model. In the top two rows, Lotka-Volterra dynamics of an autonomous model and its flow-kick counterpart are shown at three disturbance levels. In the bottom row, the invasion graph of the autonomous system (which equals the invasion graph of its flow-kick counterpart) is displayed, with nodes representing different species assemblages and arrows indicating possible transitions between communities through invasion processes. Parameter values are as in Figure 1A.
5.1. Invasion graphs for impulsive Lotka-Volterra systems
Consider the Lotka-Volterra models with impulsive disturbances introduced in Section 2.1. For these models, Section 3.2 showed that the invasion growth rate $r_{i}(\mu)$ is equal to $\left(\sum_{j}a_{ij}x_{j}^{*}(\mu)+r_{i}\right)\tau+\log(1-d_{i})$ where, generically, the mean densities $x_{i}^{*}(\mu)$ correspond to the unique solution of a system of linear equations (16). Due to this uniqueness, assumption A6 holds generically. Small perturbations of the model parameters $a_{ij},r_{i},d_{i}$ ensure that the average per-capita growth rates $r_{i}(\mu)$ are non-zero for missing species $i\notin I(\mu)$, satisfying assumption A7.
The solutions of the linear equations (16) always correspond to non-negative equilibria of the autonomous Lotka-Volterra equations:
(24) $\frac{dx_{i}}{dt}=x_{i}\left(\sum_{j}\tau a_{ij}x_{j}+\tau r_{i}+\ln(1-d_{i})\right)$
where the $\ln(1-d_{i})$ terms account for the average effects of the kicks on the long-term dynamics. Moreover, the invasion growth rates of the flow-kick system correspond to the invasion growth rates of this autonomous system.
Due to this mathematical relationship, the invasion graph $\mathcal{G}_{\mbox{\tiny flow-kick}}$ for the flow-kick system is a subgraph of the invasion graph $\mathcal{G}_{\mbox{\tiny auto}}$ for the autonomous system (24). However, $\mathcal{G}_{\mbox{\tiny flow-kick}}$ may lack some vertices of $\mathcal{G}_{\mbox{\tiny auto}}$ and their corresponding incoming and outgoing edges. Whether this occurs is an open question with ecological implications: if the graphs differ, it would suggest that disturbances fundamentally alter which community states are possible, beyond merely changing their stability properties. If $\mathcal{G}_{\mbox{\tiny auto}}$ is acyclic and satisfies the invasion condition, then $\mathcal{G}_{\mbox{\tiny flow-kick}}$ is acyclic and satisfies the invasion condition, and consequently, the flow-kick system is permanent.
An important class of shared vertices between $\mathcal{G}_{\mbox{\tiny auto}}$ and $\mathcal{G}_{\mbox{\tiny flow-kick}}$ corresponds to subsystems of the autonomous model that meet the permanence conditions of Theorem 4.8. Specifically, let $I\subset[k]$ correspond to a vertex of the autonomous invasion graph $\mathcal{G}_{\mbox{\tiny auto}}$ and define $\mathcal{G}_{\mbox{\tiny auto},I}$ to be the subgraph of $\mathcal{G}_{\mbox{\tiny auto}}$ with all vertices $J\subset[k]$ such that $J\subseteq I$ and all the corresponding edges. If $\mathcal{G}_{\mbox{\tiny auto},I}$ is acyclic and satisfies the invasion criterion of Theorem 4.8, then the same holds for the subgraph $\mathcal{G}_{\mbox{\tiny flow-kick},I}$ of $\mathcal{G}_{\mbox{\tiny flow-kick}}$. Therefore, the subsystem of the flow-kick system corresponding to the species in $I$ is permanent. By the Lefschetz [1926] fixed point theorem (which guarantees the existence of fixed points for certain classes of continuous maps on compact spaces), there exists a fixed point $z^{*}\in\mathcal{S}_{I}$ on the kick-to-kick map and, consequently, an ergodic measure $\mu=\delta_{z^{*}}$ with $I(\mu)=I$. Therefore, $I$ is a vertex of $\mathcal{G}_{\mbox{\tiny flow-kick}}$.
To illustrate how to utilize these results, consider the Lotka-Volterra model based on empirical data from four vascular plant species (Hypericum ellipticum, Juncus pelocarpus, Lysimachia terrestris, Dulichium arundinaceum) on the shoreline of Axe Lake, Ontario [Wilson and Keddy, 1986, Keddy and Shipley, 1989]. In the absence of disturbances, the invasion graph for this community is acyclic and satisfies the permanence condition: all four species coexist at a stable equilibrium.
Now consider disturbances that only impact species 5 (Figure 2). For each disturbance level $d_{5}$, I simulate the dynamics of the flow-kick system for $\tau=24$ (middle row of Figure 2) and the corresponding autonomous system (top row of Figure 2). For the autonomous system, I also calculated its invasion graph $\mathcal{G}_{\mbox{\tiny auto}}$ at all disturbance levels (bottom row of Figure 2) using the algorithm introduced in Hofbauer and Schreiber [2022]. For each disturbance level, these invasion graphs are acyclic, and the subgraph corresponding to every vertex is acyclic and satisfies the permanence condition. Hence, the invasion graph for the autonomous system is also the invasion graph for the flow-kick system: $\mathcal{G}_{\mbox{\tiny auto}}=\mathcal{G}_{\mbox{\tiny flow-kick}}$.
In the absence of disturbances ($d_{5}=0$), the autonomous and flow-kick systems are identical and the entire system is permanent (first column of Figure 2). Simulations suggest that coexistence occurs at a globally stable equilibrium. At intermediate disturbance levels ($d_{5}=0.5$), the invasion graphs $\mathcal{G}_{\mbox{\tiny auto}}=\mathcal{G}_{\mbox{\tiny flow-kick}}$, in conjunction with Proposition 4.10, predict an attractor that excludes species 2 (second column of Figure 2). At higher disturbances ($d_{5}=0.9$), the invasion graphs predict an attractor excluding species 2 and 3 (third column of Figure 2). At sufficiently high levels of disturbance-induced mortality rates ($d_{5}=0.95$), the invasion graphs predict an attractor that excludes species 2, 3, and 1.
Interestingly, species 5 is not the first to be excluded despite being the only species directly affected by disturbance-induced mortality. This unexpected result arises from the complex network of interactions: species 5 exerts positive direct and indirect effects on species 2 and 3. These interaction effects buffer species 5 against disturbance while simultaneously making its competitors more vulnerable to extinction. This highlights how indirect effects in ecological networks can lead to counterintuitive outcomes in response to targeted perturbations.
Figure 3. Coexistence and exclusion in periodic serial passage experiments. (A) Regions of coexistence and exclusion based on the signs of the average per-capita growth rates. (B) Competitive dynamics for the parameter combinations highlighted in (A). Parameters: $\delta=0.1$.
5.2. Coexistence and exclusion of competitors in serial passage experiments
Recall the competition model in Section 2.2 with extinction set $\mathcal{S}_{0}=\{(\delta,0,0),(0,\delta,0),(\delta,0,1),(0,\delta,1)\}$. For a Morse decomposition of $\mathcal{S}_{0}$, I break it up into two sets $M_{1}=\{(\delta,0,0),(\delta,0,1)\}$ and $M_{2}=\{(0,\delta,0),(0,\delta,1)\}$ consisting of the two periodic orbits of single species of the kick-to-kick map $\kappa$.
Section 3.3 identified the unique ergodic measures $\mu_{1},\mu_{2}$ supported on the sets $M_{1}$ and $M_{2}$, respectively. The average per-capita growth rates $r_{i}(\mu_{j})$ at these ergodic measures are given by (17). Lemma 3.1 implies $r_{1}(\mu_{1})=r_{2}(\mu_{2})=0$. By Theorem 4.3, the model is permanent if:
$\frac{1}{2}\sum_{y=0,1}\frac{\rho_{2}(y)}{\rho_{1}(y)}>1\mbox{ and }\frac{1}{2}\sum_{y=0,1}\frac{\rho_{1}(y)}{\rho_{2}(y)}>1.$
Equivalently,
(25) $\frac{1}{2}\sum_{y=0,1}\frac{\rho_{1}(y)}{\rho_{2}(y)}>1>\frac{1}{\frac{1}{2}\sum_{y=0,1}\frac{\rho_{2}(y)}{\rho_{1}(y)}}.$
Condition (25) is equivalent to the condition found in Yi and Dean [2013]. The left-hand side of (25) corresponds to the arithmetic mean of the ratios $\frac{\rho_{1}(0)}{\rho_{2}(0)}$ and $\frac{\rho_{1}(1)}{\rho_{2}(1)}$, while the right-hand side corresponds to the harmonic mean of these ratios. Thus, coexistence requires that the arithmetic mean of the relative fitnesses for each species is greater than the corresponding harmonic mean. This creates a “fitness balancing” requirement for coexistence.
When one of the inequalities in (25) is violated, Proposition 4.5 implies that one of the periodic orbits in the extinction set is an attractor and, consequently, the system is not permanent. For example, if
(26) $\frac{1}{2}\sum_{y=0,1}\frac{\rho_{1}(y)}{\rho_{2}(y)}<1$
then $r_{1}(\mu_{2})<0$ and $M_{2}$ is an attractor. Moreover, according to Jensen’s inequality, if (26) holds, then
$\frac{1}{\frac{1}{2}\sum_{y=0,1}\frac{\rho_{2}(y)}{\rho_{1}(y)}}\leq\frac{1}{\frac{1}{2}\sum_{y=0,1}\frac{\rho_{1}(y)}{\rho_{2}(y)}}<1.$
Thus, $r_{2}(\mu_{1})>0$ and $M_{1}$ is a repeller. This suggests that species $2$ excludes species $1$. However, whether exclusion occurs for all positive initial conditions is an open question, as the attractor-repeller structure established by our theoretical results guarantees exclusion only for some set of initial conditions but does not necessarily determine the full basin of attraction for the exclusion attractor. Using a similar argument for the reverse of the second inequality in (25) implies that, generically, the species coexist or one excludes the other. Notably, there is no bi-stability in the system. Figure 3A illustrates the main conclusions of this analysis.
Figure 4. Coexistence and exclusion in quasi-periodically forced serial passage experiments. (A) Regions of coexistence and exclusion based on the signs of the average per-capita growth rates. (B) Competitive dynamics for the parameter combinations highlighted in (A). Parameters: $\rho_{1}(\cos\theta,\sin\theta)=1.1+\Delta+\sin(\theta)$ where $\Delta$ is the mean fitness difference, $\rho_{2}(\cos\theta,\sin\theta)=1.1+\sin(\theta+\phi)$ and $\psi=1$. See the end of Section 5.2 for the modified model description.
All of these arguments extend to more complex environmental dynamics as characterized by the auxiliary variable $y$. For example, one can replace the two-periodic kick dynamics in $\{0,1\}$ with a quasi-periodic motion on the circle. Specifically, let $S^{1}=\{(\cos\theta,\sin\theta):\theta\in\mathbb{R}\}$ be the unit circle in $\mathbb{R}^{2}$. Let $\rho_{1},\rho_{2}:S^{1}\to(0,\infty)$ be two continuous positive functions. As before, define the flow by $\frac{dx_{i}}{dt}=\rho_{i}(y)x_{i},\frac{dy}{dt}=0$. The time to kick $\tau(x,y)$ is implicitly given by $\sum_{i}x_{i}\exp(\rho_{i}(y)\tau)=1/\delta$. The kick map is $H(x_{1},x_{2},\cos\theta,\sin\theta)=(\delta x_{1},\delta x_{2},\cos(\theta+\psi),\sin(\theta+\psi))$, where $\psi/\pi$ is irrational. In this case, the average per-capita growth rates in the extinction set are:
$\left(\ln\frac{1}{\delta}\right)\int_{0}^{2\pi}\left(\frac{\rho_{i}(\cos\theta,\sin\theta)}{\rho_{j}(\cos\theta,\sin\theta)}-1\right)\frac{d\theta}{2\pi}\mbox{ with }i=1,2\mbox{ and }j=1,2$
For the quasi-periodic case illustrated in Figure 4, we set $\rho_{1}(\cos\theta,\sin\theta)=1.1+\Delta+\sin(\theta)$ where $\Delta$ represents the mean fitness difference, and $\rho_{2}(\cos\theta,\sin\theta)=1.1+\sin(\theta+\phi)$ with $\psi=1$ determining the rotation angle in the kick map. These functions create environments that vary smoothly along the circle, with each species having different optimal phases.
5.3. Persistence and Extinction of the Two-Patch Pulsed Consumer-Resource Model
Figure 5. Impacts of reproductive delays and natal dispersal on consumer persistence. (A) The average per-capita growth rate $r(\mu)$ at the extinction set as a function of the reproductive delay $\tau$ for different dispersal probabilities. (B)-(D) The consumer dynamics for the reproductive lags $\tau$ and dispersal probabilities $d$ indicated in (A). Parameters: $\alpha=(0,4)$, $\beta=(0.1,0.1)$, $a=(0.1,0.1)$, $m=(0.1,0.1)$, $\theta=0.5$, $\tau$ as shown, and $d=0.0,0.125,0.25,0.375,0.5$.
Consider the model introduced in Section 2.3 with a consumer living in two patches. For this model, Section 3.4 showed that there was a unique ergodic measure $\mu$ on the extinction set $\mathcal{S}_{0}$. This invariant measure corresponds to the consumer-free equilibrium $z^{*}=(0,\alpha_{1}/\beta_{1},\alpha_{2}/\beta_{2},0,0,y_{5}^{*})$ where $y_{5}^{*}$, given by (18), corresponds to the stable patch distribution of an infinitesimally small consumer population at the consumer-free equilibrium. The average per-capita growth rate $r(\mu)$ of the consumer at this equilibrium is given by (19).
By Theorem 4.3 with the trivial Morse decomposition $M_{1}=\{z^{*}\}$ of $\mathcal{S}_{0}$, $r(\mu)>0$ implies that the consumer persists in the sense of permanence. Alternatively, if $r(\mu)<0$, then Proposition 4.5 implies that the extinction set is an attractor. Furthermore, since the resource density in patch $i$ is always $\leq\alpha_{i}/\beta_{i}$ on the global attractor $\Gamma$, $r(z)\leq r(\mu)<0$ for any $z\in\Gamma$. Consequently, the consumer-free equilibrium is globally stable whenever $r(\mu)<0$.
To explore some biological implications of this persistence condition, I consider how $r(\mu)$ depends on the dispersal probability $d$ and the reproductive delay $\tau$. As $r(\mu)$ corresponds to the dominant eigenvalue of the matrix (20), the reduction principle [Feldman and Liberman, 1986, Kirkland et al., 2006, Altenberg et al., 2017]–which states that in heterogeneous environments, random dispersal reduces population growth rates–implies that $r(\mu)$ is a decreasing function of the dispersal rate $d$ whenever $\gamma_{1}\neq\gamma_{2}$ where $\gamma_{i}=e^{-m_{i}\tau}(1+\theta\tau a_{i}\alpha_{i}/\beta_{i})$, that is, whenever there is spatial heterogeneity. In the limits of no dispersal and complete dispersal, we have:
$\lim_{d\downarrow 0}r(\mu)=\max_{i}\ln\gamma_{i}\mbox{ and }\lim_{d\uparrow 1}r(\mu)=\ln\frac{\gamma_{1}+\gamma_{2}}{2}.$
Therefore, whenever $\frac{\gamma_{1}+\gamma_{2}}{2}<1<\max_{i}\gamma_{i}$, there is a critical dispersal rate $d^{*}$ below which the consumer persists and above which the consumer goes extinct.
Alternatively, to understand the impact of reproductive delay $\tau$ on consumer persistence, I explore the behavior of $r(\mu)$ near $\tau=0$ and in the limit $\tau\to\infty$. For $\tau=0$, the matrix (20) is the doubly stochastic matrix $\begin{pmatrix}1-d&d\\ d&1-d\end{pmatrix}$ with dominant eigenvalue $1$, dominant right eigenvector $v=\begin{pmatrix}\frac{1}{2}\\ \frac{1}{2}\end{pmatrix}$, and dominant left eigenvector $w=\begin{pmatrix}1&1\end{pmatrix}$. Let $\lambda(\tau)$ be the leading eigenvalue of the matrix $A(\tau)$ given by (20). Horn and Johnson [1994, Theorem 6.3.12] implies that:
$\lambda^{\prime}(0)=wA^{\prime}(0)v=\frac{1}{2}\sum_{i}(\theta a_{i}\alpha_{i}/\beta_{i}-m_{i}).$
In particular, $\lambda(\tau)$ increases for small reproductive delays if $\frac{1}{2}\sum_{i}(\theta a_{i}\alpha_{i}/\beta_{i}-m_{i})>0$. Alternatively, $\lim_{\tau\to\infty}\lambda(\tau)=0$. Hence, if $\frac{1}{2}\sum_{i}\theta a_{i}\alpha_{i}/\beta_{i}>\frac{1}{2}\sum_{i}m_{i}$ then there are two critical values of $\tau$, $0<\tau_{*}\leq\tau^{*}$, such that $r(\mu)>0$ for $0<\tau<\tau_{*}$ and $r(\mu)<0$ for $\tau>\tau^{*}$. I conjecture that $\tau_{*}=\tau^{*}$.
Figure 5 illustrates the simultaneous effects of reproductive delay $\tau$ and dispersal probability $d$ on consumer persistence. In this figure, patch $1$ has no resource production ($\alpha_{1}=0$) and therefore does not support consumer reproduction, while patch $2$ has positive resource production ($\alpha_{2}=4$) sufficient to support consumer reproduction. As $\frac{1}{2}\sum_{i}\theta a_{i}\alpha_{i}/\beta_{i}>\frac{1}{2}\sum_{i}m_{i}$, $r(\mu)$ is positive for sufficiently small reproductive delays and negative for sufficiently long reproductive delays (Figure 5A). Consistent with my conjecture that $\tau_{*}=\tau^{*}$, the low-density consumer growth rate $r(\mu)$ has a unique maximum with respect to $\tau$. Moreover, consistent with the reduction principle, $r(\mu)$ decreases with the probability of dispersal. Hence, intermediate reproductive delays or lower dispersal probabilities are required for consumer persistence (Figure 5B–D).
6. Discussion
Natural populations experience a complex interplay of continuous and discrete processes: continuous growth and interactions are punctuated by discrete reproduction events, dispersal, and external disturbances. This work developed a mathematical framework that characterizes permanence in flow-kick systems through invasion growth rates. By accommodating auxiliary variables and the state-dependent timing of discrete events, the framework extends beyond previous work on purely continuous or discrete systems [Roth et al., 2017, Patel and Schreiber, 2018, Hofbauer and Schreiber, 2022]. However, several mathematical challenges and opportunities for improvement remain. Applying these results to three ecological models reveals how the interplay between continuous and discrete dynamics can lead to some counterintuitive predictions for population persistence and ecological outcomes.
6.1. Mathematical Advances and Challenges
The permanence results in this paper involved defining the invasion growth rates and proving the permanence of the discrete-time kick-to-kick map $\kappa$. The assumptions of dissipativeness, positivity and continuity allowed the extension of permanence from the discrete-time map to the continuous-time flow-kick system (Lemma 4.1).
Theorem 4.3 extends the work of Roth et al. [2017] of discrete-time maps in three important ways. First, it generalizes their framework through the use of auxiliary variables. This generalization is significant because, although Roth et al. [2017] developed invasion growth rate characterizations of permanence for discrete-time maps with stage structure, our approach encompasses a much broader class of systems. As illustrated by the spatially structured consumer model in Section 2.3, stage structure can be incorporated by a change of variables corresponding to the total population density (i.e., $x_{i}$) and the frequencies in the different stages (that is, part of the variables in the auxiliary state $y$). However, our auxiliary variables can also represent deterministic forcing (e.g. periodic, quasiperiodic, chaotic) of the community dynamics as illustrated in Section 5.2, trait evolution [Patel and Schreiber, 2018], plant-soil feedbacks, and many other processes. Moreover, by using auxiliary variables, our theorem makes no assumptions about the irreducibility of stage structure models, in contrast to Roth et al. [2017]. Building on this first advancement, the second contribution is methodological: our proof of Theorem 4.3 is substantially simpler than the approach in Roth et al. [2017]. Their method required extending the models to skew product systems in an infinite-dimensional sequence space (i.e., the inverse limit of the system) and relied on non-autonomous Perron-Frobenius theorems, such as those in Ruelle [1979]. Our approach avoids these complications. Finally, from a theoretical perspective, Theorem 4.8 provides the first proof that uses invasion graphs to characterize permanence for discrete-time models.
The key lemma for the proofs (lemma 4.4) simplifies the earlier proofs using ergodic theory to characterize permanence [Schreiber, 2000, Hofbauer and Schreiber, 2010, Roth et al., 2017, Hofbauer and Schreiber, 2022] and provides a unified approach to the two permanence theorems and the different model types. In forward time, this lemma has a simple biological interpretation. If all species are initially present, then the invasion growth rates associated with the community trajectory are non-positive as their densities remain bounded. Conversely, going backward in time, (bounded) trajectories are associated with non-negative invasion growth rates. Namely, they were either rare in the past and became more common (i.e., a positive invasion growth rate) or they were common and remained common (i.e. a zero invasion growth rate).
Despite this progress, there are multiple mathematical opportunities and open challenges. Most importantly, in the case of cyclic invasion graphs, there is no general criterion for permanence. Theorem 4.3 suggests that requiring the Hofbauer condition (Equation 23) for each connected component of the invasion graph should be sufficient (see, e.g., the conjecture in Spaak et al. [2023]). However, this has yet to be proved. Even if such a result holds, a substantial gap will remain. Specifically, there is a gap between the sufficient conditions for permanence in Theorem 4.3 and the sufficient condition in Proposition 4.5 for the existence of extinction-bound trajectories. This gap arises due to the potential formation of complex networks of heteroclinic orbits (trajectories connecting different equilibrium points) in ecological models [Hofbauer, 1994, Schreiber and Rittenhouse, 2004, Hofbauer and Schreiber, 2004, Hofbauer, 2007, Spaak et al., 2023].
Beyond these finite-dimensional models, there is a need to develop a more comprehensive approach based on invasion growth rates for infinite-dimensional impulsive systems. These models are widely used in mathematical ecology and epidemiology [Luo and Wang, 2021, Fazly et al., 2017, Yang and Zhong, 2014, Lewis and Li, 2012]. Part of the challenge is the development of the appropriate spectral theory for non-autonomous operators. However, I speculate that the use of auxiliary variables may provide a useful way to circumvent some of these challenges.
6.2. Biological Insights and Opportunities
Applying the invasion growth rate criterion to three ecological models provides mathematically rigorous extensions of previously observed phenomena, offers several new biological insights, and highlights opportunities for future research.
The analysis of Yi and Dean [2013]’s model of microbial serial transfer experiments demonstrates how state-dependent timing of kicks (dilution to new containers) can act as a coexistence mechanism. In this model, competing species grow exponentially between dilution events in environments that alternately favor one species or the other. As observed by Yi and Dean [2013], if dilutions occur at regularly spaced time intervals, competitive exclusion results: the species with the highest average per-capita growth rate excludes the other species (i.e., $r_{i}(\mu)=\tau(\sum_{y}\rho_{i}(y))$).
In contrast, when the timing of the dilution depends on the critical total population size ($x_{1}+x_{2}$), coexistence becomes possible through a storage effect mechanism [Li and Chesson, 2016, page E55]. This occurs because faster-growing species in any given environment reach higher densities more quickly, triggering earlier dilution events. These earlier dilutions effectively reduce the competitive advantage of the temporarily favored species, allowing both species to persist. This represents a classic fluctuation-dependent coexistence mechanism [Chesson, 1994].
The analysis presented here demonstrates that these conclusions hold even when the environment exhibits greater complexity in environmental fluctuations, e.g., with more than two environmental states or with quasiperiodic environments. More importantly, my analysis suggests that there are only two possible ecological outcomes: coexistence or exclusion. In particular, bistability, as observed in classical Lotka-Volterra competition models, is not possible, a conclusion not observed by Yi and Dean [2013]. However, verifying that exclusion occurs for all positive initial conditions remains an open mathematical challenge.
The analysis of a two-patch version of Pachepsky et al. [2008]’s consumer-resource model reveals a counterintuitive effect of reproductive delays on consumer persistence. Pachepsky et al. [2008] likely didn’t comment on this effect due to their non-dimensionalization of time in the model. To understand this phenomenon, first consider the simpler case of a spatially homogeneous (or single patch) model. Here, the consumer invasion growth rate depends on two competing factors. First, the probability of survival to reproduction ($e^{-m\tau}$) decreases exponentially with longer delays ($\tau$). Second, reproductive output ($1+\theta\tau a\alpha/\beta$) increases only linearly with longer delays. This creates a fundamental trade-off: longer reproductive delays allow more time for offspring production, but the survival probability decreases more rapidly than offspring production increases.
As a result of this trade-off, when the rate $\theta a\alpha/\beta$ of potential offspring production exceeds the per-capita mortality rate $m$, the per-capita growth rate is maximized at an intermediate reproductive delay of $\frac{1}{m}-\frac{1}{\theta a\alpha/\beta}$. The role of spatial heterogeneity in optimal reproductive delay is an interesting question that deserves further analysis.
Finally, the analysis of the empirically parameterized Lotka-Volterra model reveals a counterintuitive relationship between facilitation and disturbance effects on community composition. When competitors facilitate each other’s growth (directly or indirectly), disturbance targeting one species can cascade through these positive interactions, potentially causing extinction of species not directly affected by the disturbance. This presents a paradox: the very facilitative interactions that might promote diversity under stable conditions can amplify vulnerability when disturbances occur. Given increasing empirical evidence that facilitation is more common in plant communities than previously thought [Burns and Strauss, 2011, Siefert et al., 2018, Buche et al., 2025], using the methods presented here to explore the interaction between facilitation and disturbance in community composition provides exciting opportunities for future research.
On the more mathematical side, our analysis of the Lotka-Volterra systems showed that the invasion graph for a flow-kick system is always a subgraph of the invasion graph for an appropriately defined autonomous Lotka-Volterra model. Understanding whether there are cases where it is a proper subgraph remains an interesting yet challenging mathematical problem.
6.3. Concluding remarks
In conclusion, this work advances both the mathematical theory of permanence and our ecological understanding of how continuous and discrete processes interact to influence species persistence. The invasion growth rate framework developed here provides a powerful tool for analyzing coexistence and extinction that complements existing approaches in theoretical ecology. Although important mathematical challenges remain—-particularly for systems with cyclic invasion graphs—the applications presented demonstrate the framework’s ability to reveal counterintuitive ecological dynamics. These range from how state-dependent disturbances can promote coexistence in microbial communities to how facilitation networks can amplify disturbance effects in plant communities.
Future work that incorporates stochasticity, develops computational methods to calculate invasion graphs in a mathematically rigorous manner, and expands applications to new ecological contexts will further enhance our ability to predict and manage the dynamics of natural populations experiencing both continuous growth and discrete disturbances.
Acknowledgments
This work was partially funded by National Science Foundation Grant DEB-2243076. The author thanks Jakob Karre-Rasmussen for constructive feedback on earlier versions of the manuscript and the participants of the joint Hastings-Schreiber Lab Meetings in Spring 2024 for inspiring work on flow-kick dynamics.